OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that satisfies the following condition. The line passes through (25, 3), and is parallel to the graph of y =9/4x + 5. If a decimal, round to the nearest whole number

OpenStudy (anonymous):

i need ans

OpenStudy (anonymous):

i aint got noone to help me with this no teacher just computer stuff

OpenStudy (anonymous):

?????

OpenStudy (jack1):

y = mx + c is the slope intercept form

OpenStudy (jack1):

parallel to the graph of y =9/4x + 5, so has the same gradient (m) value

OpenStudy (jack1):

and use one of the points it passes through to find the y intercept (c)

OpenStudy (anonymous):

i need answer

OpenStudy (jack1):

so work it out do the math yourself and i'll help you, but i'm not an answer generator dude

OpenStudy (jack1):

would you like more clarification on any of the points...?

OpenStudy (anonymous):

na im good

OpenStudy (anonymous):

@Jack1 can you help me on this problem as well,

OpenStudy (anonymous):

slope intercept i though it was y=mx+B

OpenStudy (jack1):

no worries the final equation should be in the slope intercept form, so y = mx + c or y = mx + b if that's the way you were taught, either works now we need to work out "m" and "b" so...

OpenStudy (jack1):

m is the gradient so if a line is parallel to our line, then it must have the same gradient so we have a parallel line: \[y =(\frac{ 9 }{ 4 }) x + 5\] so first step, steal the gradient from there now all we have left is the B value ( y intercept)

OpenStudy (jack1):

to find that, we need a point along our line we're given the point (25, 3) hang on... are you guys following ok...?

OpenStudy (anonymous):

yes

OpenStudy (jack1):

cool, so what's your equation so far? y = mx + B so your value for m is...?@amandasoto144

OpenStudy (anonymous):

25?

OpenStudy (jack1):

no... 9/4, same as the other given equation

OpenStudy (anonymous):

oh i didnt see that part till now sorry

OpenStudy (anonymous):

so b is 5?

OpenStudy (jack1):

all good ok, so we have the point (25, 3) somewhere on our line (which is an x, y point) and we have the gradient so if y = mx + B we can use 3 = 9/4 * 25 + B to find b

OpenStudy (jack1):

no, only the gradients are the same in parallel lines, not the y intercepts (b values)

OpenStudy (jack1):

so 3 = 56.25 + b therefore b = ...?

OpenStudy (anonymous):

im sorry im confused, so i know how you got 56.25 but am i suppose to add 5 or 3?? how am i suppose to find B?? or is 56.25 B??

OpenStudy (anonymous):

is the answer 59.25??

OpenStudy (jack1):

no, we dont know B yet, so we need to isolate it in the equation y = mx + b 3 = 56.25 + b ( so take 56.25 from both sides) 3 - 56.25 = 56.25 + b - 56.25 -53.25 = b

OpenStudy (jack1):

so your final equation is y = m x + c just sub in the values of m and b you worked out

OpenStudy (jack1):

* b or * c... whatever you want to call it, sorry

OpenStudy (anonymous):

so Y=9/4x+53.25

OpenStudy (anonymous):

?

OpenStudy (jack1):

almost, negative value for b remember

OpenStudy (jack1):

-53.25

OpenStudy (anonymous):

but you said plug it in from both sides, and i saw two negatives so i thought it was positive, my fault sorry

OpenStudy (anonymous):

so Y=9/4x-53.25

OpenStudy (jack1):

perfect

OpenStudy (anonymous):

oh wow, thank you!

OpenStudy (jack1):

s'welcome